OpenStudy (anonymous):

Which of the following is equivalent to three and two-fifths divided by four? A.one and three-seventeenths B.three and one-tenth C.fifteen twentieths D.seventeen-twentieths

OpenStudy (texaschic101):

do you know how to divide fractions ?

OpenStudy (goldphenoix):

\[3\frac{ 2 }{ 5 } \div 4\] To find the answer you can use KCF (Keep, Change, Flip) But before you do that, you must convert the mixed number into impropre fraction. To do this, time the denonimator by the whole number and add it by 2. Keep the denonimator. So it would be:\[\frac{ 17 }{ 5 } \div 4\] Now use KCF. You keep the first fraction. You change the division symbol into a multiplication symbol. You flip 4 into 1/4. So it would finally look like: \[\frac{ 17 }{ 5 } \times \frac{ 1 }{ 4} \] Now multiply the fraction regularly. 17 x 1 =17. 5x4 = 20 Your answer should be: \[\frac{ 17 }{ 20 }\]

OpenStudy (goldphenoix):

So the answer would be D. I hope this help. :)

OpenStudy (texaschic101):

good explaining @GoldPhenoix ....very understandable :)

OpenStudy (goldphenoix):

Thank you. :)